OpenStudy (anonymous):

Can y'all help me with the rest of these to make sure im doing it right... 3^6/3^2 , 9^0 ,(5^3)^2 , (-4)^4. Simplify and evaluate :)

OpenStudy (johnweldon1993):

When dividing exponents with the same base (3^6/3^2) you leave the base (3) alone and just subtract the exponents... so (3^6/3^2) = 3^(6 - 2) = 3^4 and evaluate that...

OpenStudy (johnweldon1993):

9^0 general rule...what does any number raised to the 0 power = ?

OpenStudy (johnweldon1993):

(5^3)^2 when exponents are separated by parenthesis like this...you multiply the exponents...so (5^3)^2 = 5^(3 * 2) = 5^6 and evaluate that...

OpenStudy (johnweldon1993):

And finally (-4)^4 just multiply -4 x -4 x -4 x -4 and you get...

OpenStudy (anonymous):

Thank you ! I was doing it right all along :)

OpenStudy (johnweldon1993):

Anytime! and Great work :)